OpenStudy (anonymous):

Whole under root of 2+ 2 under root 3 iota?

jhonyy9 (jhonyy9):

???

OpenStudy (anonymous):

you didnt get the question?

jhonyy9 (jhonyy9):

yes

jhonyy9 (jhonyy9):

i dont understand your question

OpenStudy (anonymous):

square root of 2+2(square root 3 )i \[\sqrt{2+2\sqrt{3}}\]i

jhonyy9 (jhonyy9):

ok what you need to do with ?

OpenStudy (anonymous):

we need to find the square root

OpenStudy (anonymous):

and give the final answer in form of iota

jhonyy9 (jhonyy9):

what is this ,,i" ? or this wann being a complex number in the form of a +bi ?

OpenStudy (anonymous):

yes its complex number. i is iota

OpenStudy (unklerhaukus):

what to do with this expression?

OpenStudy (anonymous):

you have to find the square root of the complex number

OpenStudy (unklerhaukus):

\[\sqrt{2+2\sqrt3 i}\]

OpenStudy (anonymous):

yes